Grade 11Mechanics

The kinetic energy K of a particle moving along a circle of radius(R)depends on the distance covered (s) as K = cs^2.The force acting on the particle is

Blazing Rig
7 Years agoGrade 11

1 Answer

Khimraj
7 Years ago
K = cs2
½*mv2 = cs2
differentiating wrt time
mva = 2csv
ma = 2cs
Tangential force
Ft = ma = 2cs
centripetal force = mv2/R =  2cs2/R
Total force F = \sqrt{F_{t}^{2} + F_{c}^{2}}
SO F = 2cs\sqrt{1 + s^{2}/R^{2}}
Hope it clears.
If you like answer then please approve it.